OpenStudy (anonymous):

is y = cos^-1(x) SAME AS x = cos(y), where 0 < y < π

OpenStudy (anonymous):

or x = 1/cos(y), where y is any number other than π/2 times an odd number

OpenStudy (anonymous):

go with the first one, since arccosine is a well defined function

OpenStudy (anonymous):

how about this one y = 1/cos(x), where x is any number other than π/2 times an odd number

OpenStudy (anonymous):

wouldnt y=1/cos(x) be the same as y = cos^-1(x)